OpenStudy (anonymous):

(e^x)^x*e^24=e^12x solve

OpenStudy (carson889):

(e^x)^x = e^(x^2) e^(x^2) * e^24 = e^(x^2 + 24) Therefore, e^(x^2 + 24) = e^(12x) Take the ln of both sides: \[\ln(e ^{x ^{2}+24}) =\ln(e ^{12x}) \rightarrow x ^{2} + 24 = 12x \rightarrow x ^{2} -12x + 24 = 0\] Use quadratic equation: (-b +- sqrt(b^2 - 4ac))/2a (-(-12) +/- sqrt(144 - 4(1)(24))/2 = (12 + sqrt(144-96))/2 = 6+sqrt(48)/2 = 6 + 2sqrt(3) Other x = 6 - 2sqrt(3)